OpenStudy (vanz):

2x+y=14 plug in 3/5 for x ??

OpenStudy (rulnick):

2(3/5)+y=14 6/5+y=14 y=14-6/5 y=12 4/5 or y=64/5.

OpenStudy (vanz):

if i were to get ride of the fraction 3/5 by multiplying everything by 5, how would that look??/

OpenStudy (vanz):

???????

OpenStudy (anonymous):

it would look just fine.

OpenStudy (anonymous):

except you'll just have to divide everything by 5 again in order to isolate y. so it would be kind of a pointless intermediate step

OpenStudy (vanz):

when you multiply everything by and you have for example 2(3/5)+y=14, 10(3)+5y=70, would you add or multiply the 3

OpenStudy (anonymous):

y=11.7

OpenStudy (anonymous):

2x.6=1.2 1.2y=14 so divide both sides by 1.2 to isolate Y and you get Y=11.7